### v1

# Empirical Dataset for Image Transformations

Abstract
Existing image editing models struggle to meet real-world demands; despite excelling in academic benchmarks, we are yet to see them adopted to solve real user needs. The datasets that power these models use artificial edits, lacking the scale and ecological validity necessary to address the true diversity of user requests. In response, we introduce RealEdit, a large-scale image editing dataset with authentic user requests and human-made edits sourced from Reddit. RealEdit contains a test set of $9.3$K examples the community can use to evaluate models on real user requests. Our results show that existing models fall short on these tasks, implying a need for realistic training data. So, we introduce $48$K training examples, with which we train our RealEdit model. Our model achieves substantial gains—outperforming competitors by up to $165$ Elo points in human judgment and $92\%$ relative improvement on the automated VIEScore metric on our test set. We deploy our model back on Reddit, testing it on new requests, and receive positive feedback. Beyond image editing, we explore RealEdit ’s potential in detecting edited images by partnering with a deepfake detection non-profit. Finetuning their model on RealEdit data improves its F1-score by $14$ percentage points, underscoring the dataset’s value for broad, impactful applications.
Figure 1: We visualize edits made by our model. We introduce RealEdit, a large-scale image editing dataset sourced from Reddit with real-world user edit requests and human-edits. By finetuning on RealEdit, our resultant model outperforms existing models by up to 165 Elo points with human judgment and delivers real world utility to real user requests online.
1 Introduction
The need to edit photos is more important than ever—people everywhere seek to perfect, enhance, or restore their images, from casual snapshots to treasured memories. If more effective and aligned editing models were readily available, many would use them for a variety of purposes: to remove an unwanted photobomber, adjust lighting in their selfies, restore their grandparents’ wedding photos, or even add creative effects. This demand is vividly demonstrated in online communities like Reddit’s r/PhotoshopRequest [^1] https://www.reddit.com/r/PhotoshopRequest and r/estoration [^2] https://www.reddit.com/r/estoration, with over 1.5 million combined members. Many users pay money for quality edits, highlighting the demand for advanced, user-friendly editing tools.
Despite the impressive capabilities in image generation and modification led by recent advancement of diffusion models [45, 46, 6, 65], seemingly straightforward real-world editing tasks, like ones from the Reddit’s r/PhotoshopRequest, continue to pose significant challenges to existing models. For instance, while existing models are effective at artistic transformations or generating stylized content [65, 36, 28, 35, 48], they fall short at some of the most common real-world requests such as restoring a damaged image (see Figure 2). This discrepancy highlights a critical misalignment between the capabilities of current editing models and the actual needs of users.
One major challenge for models to effectively tackle real-world image editing is the diversity and open-ended nature of the tasks involved. However, most existing models are trained with synthetic or arbitrarily created datasets that do not characterize human-centered objectives well, as is shown in Table 1. For example, in Ultra-Edit [69], “adding a rainbow” to an image constitutes a significant portion of the data set. As a result, models trained on these datasets struggle to address the practical needs of real-world users.
In this work, we introduce RealEdit, a large-scale text-guided image editing dataset meticulously compiled from Reddit. RealEdit, by design, more faithfully reflects the distribution of image editing needs. Specifically, we source image editing requests from two of the largest relevant subreddit communities, r/PhotoshopRequest and r/estoration, into a dataset consisting of over $57$K editing examples, wherein each example comprises of an input image, an instruction, and one or multiple edits performed by humans. Overall, there are a total of $151$K input and edited images in this collection. By carefully preprocessing and filtering out ambiguous and noisy examples with meticulous manual verification, we transform part of the collected examples in RealEdit into an evaluation set that consists of more than $9.3$K real-world image editing requests to test models’ capability. Notably, RealEdit evaluation set shows that real-world requests differ drastically from existing evaluation datasets [65, 48], on which existing models struggle.
To build an effective image editing model for real-world tasks, we finetune a new text-guided image editing model, on RealEdit’s training examples. To produce useful edits that preserve the identities of the people in photos, we upgrade InstructPix2Pix [6] by replacing its Stable Diffusion [46] decoder with OpenAI’s Consistency decoder [40], which was pretrained on more human-centric data.
Figure 2: Baselines struggle on simple, practical tasks, such as restoring a damaged photograph. Our model is successful.
Our model demonstrates significantly better performance than existing state-of-the-art models on RealEdit’s test set with a human preference (N=4,196) Elo score of $1184$, beating the next best model by $165$ points. We also outperform existing models using automated metrics: our model achieves $4.61$ VIEScore versus the next best score of $2.4$ (amongst other metrics). Moreover, our model still remains competitive with MagicBrush and EmuEdit [65, 48] on their test set. We further validate our model by completing new Reddit requests and receiving positive feedback.
Finally, we partner with <REDACTED>, a non-profit aimed at AI-generated content detection. By adding human-made edits from our dataset, we improve their model’s F1-score by $14$ points. Our ecologically valid experimentation highlights the dataset’s value outside of editing tasks.
2 Related work
Image editing datasets. While extensive datasets exist for captioning and identifying edited images within fixed domains [10, 41], there is a notable lack of large-scale, human-edited image datasets. Currently, larger-scale image editing datasets mostly rely on synthetic data [6, 65, 48, 66, 69], while the ones with human edited images are limited in size [49, 50]. While synthetic datasets may include human inputs, such as generating instructions or ranking edits [65, 66, 6], these datasets do not contain edits that are completed by humans. Most importantly, existing datasets are curated in ways that do not necessarily characterize real-world editing distribution well. We compare RealEdit to existing datasets in Table 1.
Text-guided image editing.
There is a rich literature in models focusing on specific image editing tasks, such as inpainting [62], denoising [18], and style transfer [15]. Recent advancements emphasize generalized models that better align with human use cases, leading to innovative methods such as generating programs to modify images [19], as well as end-to-end diffusion-based or GAN-based editing models [25, 43, 2, 35, 54, 58]. Diffusion models like Stable Diffusion [46] excel at generating images from text prompts, serving as versatile models for image generation [63]. Several models [6, 36, 27] utilize diffusion-based techniques for editing, though generating images from captions alone may compromise fidelity. To mitigate this, some models [6, 36, 65, 28] leverage Prompt-to-Prompt technique [20], employing cross-attention maps to preserve most of the original image. Others achieve consistency by fine-tuning diffusion models to reconstruct images using optimized text embeddings, blending these with target text embeddings [27]. However, limitations persist, such as stuggles with face generation [5] and cross-attention requiring minimal, often single-token caption variation.
Evaluating image editing models.
Originating from early text summarization in NLP [37], QA-based evaluation methods automatically transform prompts into questions and use them to validate generated content [13, 11, 14]. In text-to-image generation, VQA-based evaluation methods transfer text into atomic questions and conduct VQA to verify generated images, providing enhanced fine-grained and interpretable benchmark results [8, 32, 7]. Notably, TIFA [22] pioneered the use of VQA for automatic evaluation, while subsequent works enhanced model-human correlation [60, 34], incorporated additional modules and MLLM-as-a-Judge [16, 9, 61, 7, 29]. To evaluate image editing models, we follow and extend existing work [48] in casting the evaluation into image generation evaluation wherein we measure the faithfulness of the edited images to their target output captions, using the aforementioned VQA-based frameworks.
3 RealEdit
We introduce RealEdit: a high-quality large-scale dataset for text-guided image editing. RealEdit dataset includes 48K training data points and 9K test data points, each featuring an original image, an editing instruction, and one to five human-edited output images. Altogether, we are publishing a total of 151K images. RealEdit is the first large-scale image editing dataset wherein real-world users both submit and complete the requests (Table 1).
Table 1: Size and human involvement across editing datasets. RealEdit is the largest dataset containing human-made edits.
Dataset Size Large Human Human Real-world
scale? select outputs? edit? requests?
InstructPix2Pix [6] 454K
MagicBrush [65] 10K
EmuEdit [48] 10M
HIVE [66] 1.1M
UltraEdit [69] 4M
AURORA [28] 280K
IER [50] 4K
GIER [49] 6K
RealEdit (Ours) 57K
Figure 3: Dataset curation pipeline. We source data from r/estoration and r/PhotoshopRequest. From the posts, we extract input images and edit instructions. The instructions are processed using a VLM to isolate the editing task. From the comments, we collect up to 5 human-edited outputs per post.
3.1 Dataset creation pipeline
The extensive and structured nature of Reddit makes it an ideal source for creating diverse large-scale datasets rooted in real-world content. We leverage this by developing a data collection pipeline with three key steps: (1) collecting raw post and comment data from the subreddits of interest, (2) processing and organizing the data, and (3) manual verification to ensure safe and high-quality outputs (Figure 3). Step 1: Subreddit selection. We build a diverse image editing dataset from two key subreddits to cover a wide range of tasks. The main source, r/PhotoshopRequest, provides 261K posts and 1.1M comments on tasks ranging from object removal and background changes, to creative edits. Additionally, we source requests from r/estoration for their sentimental value to users. This subreddit contributes 20K posts and 126K comments focused on restoring old photos, including repairing creases, colorizing black-and-white images, and enhancing clarity. We exclude larger communities like r/photoshopbattles due to their emphasis on humor and less specific editing needs. The dataset consists of original image URLs sourced from posts, edit instructions, and edited image URLs taken from the comments. The images we collected were posted between 2012 and 2021 which implies low likelihood of AI-generated content.
Step 2: Instruction refinement and caption generation. One challenge in collecting web-crawled data is that user-provided instructions may be noisy, often including personal anecdotes or task-irrelevant details (e.g., “This photo was taken of my Mother and me at my Grandmother’s wake. I would love to get this framed for my Mom’s birthday next month. I love the photo, but the person who took it put filters all over it. I was wondering if someone could make it look more natural.”). We use GPT-4o [39] to summarize the text to only the key editing requirements. The pipeline refines the noisy instruction above into the following: “Restore image damage and enhance clarity”.
For the RealEdit test set, we generated captions for both input and edited images using vision-language models to support evaluation on caption-based metrics. Implementation details are provided in the Appendix.
Step 3: Data verification and final composition. After generating the dataset, we conducted a rigorous multi-stage verification process to ensure data quality. All images were screened for inappropriate content using the opennsfw2[4] network, filtering out those flagged as explicit. Additionally, RealEdit test set was manually reviewed by two annotators evaluating the following criteria: (1) appropriateness of the input image, (2) applicability of the instruction, and (3) correctness of the output image. Approximately 78% of the data points were agreed upon as high quality and included in the final test set for RealEdit.
4 RealEdit dataset analysis
RealEdit provides insight into practical applications of image editing by analyzing real-world requests. We observe notable differences between RealEdit and existing datasets including InstructPix2Pix [6], MagicBrush [65], Emu Edit [48], HIVE [66], Ultra Edit [69], AURORA [28], Image Editing Request [50] and GIER [49]. While we focus primarily on differences with MagicBrush [65] and Emu Edit [48] in the following discussions, these observations broadly apply across datasets used to train image editing models. Figure 4 details the main differences.
Qualitative analysis and taxonomy.
We create a taxonomy of image editing tasks people have requested. This involves (1) categorizing our edit requests into operations, (2) subcategorizing requests by subject of the edit, and (3) prompting GPT-4o [39] to categorize the request based on the input image and edit instruction. To determine the operations, we modify the MagicBrush [65] set of operations for clarity. We base our possible subjects on the significant categories from the sample of 500 images. We tune our GPT-4o prompt using samples of 100 data points to ensure accuracy, then validate on a separate sample to avoid overfitting. We find that both the categories and their distribution differ greatly from prior work. Since our categorizations are fairly similar to MagicBrush [65] and Emu Edit [48] test sets, we run our taxonomy on these test sets and highlight key distributional differences in Figure 4. Full taxononomies and comparisons are listed in Appendix.
Figure 4: Key differences in the distribution of our test set compared to MagicBrush and Emu Edit test sets. MagicBrush and Emu Edit tend to be similar in distribution to each other, but starkly different from RealEdit.
Differences in edit operations.
Synthetic datasets contain a greater use of “add” requests (36% less than MagicBrush than). In contrast, real-life photos typically contain the intended objects within the frame, with many of RealEdit’s semantically focused tasks involving the removal of unintended elements, such as strangers in the background, shadows on faces, or cars on the street. Additionally, there are numerous cases where input images are semantically aligned with the owner’s intent, but errors in photography such as bad lighting, motion blur, or graininess. Following this, RealEdit contains more “enhance” requests (14% greater than MagicBrush and Emu Edit) compared to existing datasets. These findings indicate that real users often prioritize subtler requests, whereas synthetic datasets are dominated by larger semantic changes, such as “add.”
Differences in image content.
Analysis on 500 samples reveals that around 55% of the input images feature people as the main subject. Consequently, the subjects of the requested edits are more likely to be people (13% more than Emu Edit), and less likely to be man-made objects (20% less than MagicBrush). Animals and media (characters, movie/book posters, memes, etc.) are the next most common categories, comprising about 10% of the test set each. Common media requests include restoring old photographs, participating in fandoms, making memes, or other forms of online entertainment. The fixation on media is not paralleled in other datasets (15% more than Emu Edit). These findings reveal a clear difference: Reddit users tend to prioritize personal significance by including people and entertainment by incorporating media, and synthetic datasets often fail to reflect these preferences accurately.
Given the substantial distributional differences of RealEdit compared to existing datasets, we demonstrate in Section 6 that current models struggle to perform well on real-world requests.
5 An editing model trained with RealEdit
To demonstrate the value of RealEdit, we develop an image editing model using training examples in RealEdit. Specifically, we utilize InstructPix2Pix [6] as the model backbone on which we finetune using our data. We leave exploration on using different base models as future work.
Table 2: Quantitative evaluation on RealEdit test set. On all metrics other than pixel distance, the RealEdit model scores the highest.
Model VIES_O $\uparrow$ VIE_PQ $\uparrow$ VIE_SC $\uparrow$ VQA_llava $\uparrow$ VQA_Flan-t5 $\uparrow$ TIFA $\uparrow$ L1 $\downarrow$ L2 $\downarrow$ CLIP-I $\uparrow$ DINO-I $\uparrow$ CLIP-T $\uparrow$
AURORA [28] 2.20 3.43 2.40 0.606 0.711 0.724 0.154 0.069 0.793 0.733 0.246
HIVE [66] 1.73 3.40 1.86 0.596 0.678 0.685 0.246 0.142 0.743 0.646 0.250
InstructPix2Pix [6] 1.64 3.12 1.76 0.594 0.650 0.698 0.181 0.073 0.752 0.638 0.244
MagicBrush [65] 1.87 3.88 1.89 0.620 0.726 0.741 0.138 0.064 0.830 0.782 0.251
Null-text Inv. [36] 1.89 3.27 2.14 0.637 0.751 0.731 0.152 0.067 0.743 0.669 0.261
SDEdit [35] 0.59 1.47 0.75 0.588 0.653 0.703 0.156 0.068 0.678 0.613 0.230
RealEdit (Ours) 3.68 4.01 4.61 0.660 0.795 0.751 0.143 0.066 0.840 0.792 0.261
Aligning with pretraining data.
Since we finetune InstructPix2Pix rather than training the model from scratch, we align our finetuning dataset with the data distribution used in InstructPix2Pix’s pretraining data to avoid substantial distributional shifts that may deteriorate model’s performance. In particular, InstructPix2Pix [6] applies CLIP-based [44] filtering to ensure the quality of image pairs. In addition, as it employs Prompt-to-Prompt [20] in generating its training data, the input-output image pairs are with high structural similarity. To align our training set, we thus follow the same CLIP-based filtering and additionally use SSIM [55] to include structurally similar images, recognizing that human edits collected in RealEdit often alter structure with techniques like drag-and-drop adjustments and symmetrical flipping. Tasks incompatible with InstructPix2Pix’s capabilities, such as resizing images, changing file types, or highly ambiguous prompts (particularly those involving humor) are thus excluded. In total, we trained on 39K examples. Aligning our training data with the InstructPix2Pix distribution allows for more competitive performance on metrics, and accounts for limitations in the InstructPix2Pix’s architecture and pretraining. For training our model, we closely follow the configuration of MagicBrush [65]. Specifically, we train our model for 51 epochs, utilizing cosine learning rate decay and incorporating a learning rate warm-up phase (details in Appendix).
Decoding at inference.
We observe that Stable Diffusion [46] struggles with accurately reconstructing human faces and fine-grained image details.As shown in Section 4, real-world requests are human-centric with detailed edit needs. To address this, we incorporate OpenAI’s Consistency Decoder [40] at inference time, significantly enhancing generation quality for faces, patterns, and text without altering the diffusion process.
6 Experiments
Setup. We benchmark our model against six open-source baselines: InstructPix2Pix [6], MagicBrush [65], AURORA [28], SDEdit [35], HIVE [66], and Null-text Inversion [36]. We leverage the input and output captions generated in Section 3 for models that require them.
To evaluate the models, we adopt a comprehensive suite of metrics. First, we utilize VQA-based automated metrics to measure task completion, as these metrics have been shown to closely reflect human judgments. In particular, we use VIEScore [29] with a GPT-4o backbone as our default metric, as it evaluates semantic consistency (VIE_SC), perceptual quality (VIE_PQ), and overall alignment with human-requested edits (VIE_O) each on a scale of 0 to 10. Similarly, we use VQAscore [32] (with different base models: LLaVa and FLAN-T5) and TIFA [22] to evaluate the fine-grained faithfulness of the output image to the edit instruction. We also include standard metrics such as L1- and L2 pixel distance, DINO [64], CLIP-I and CLIP-T, following prior work [65, 48]. Most importantly, we leverage real users to make pairwise comparisons between edits and compute Elo ranking of the models [23]. We further qualitatively study the response Reddit users have on edits produced by our model on recent posts.
6.1 Automated evaluations on RealEdit test set
In Table 2, we show that existing models struggle to capture the semantic nuances of human requests, while our model achieves notable improvements, particularly in VIE_SC scores. Our model also significantly outperforms other baselines on finer-grained metrics like VQAScore and TIFA. Although our model achieves state-of-the-art (SOTA) results on standard metrics, these metrics are limited as they fail to fully capture task completion. Notably, using the input image as the output yields the highest scores on four out of five metrics, with the fifth, CLIP-T, exhibiting saturation effects. This underscores the importance of more nuanced automated metrics, such as VQA-based approaches, to better evaluate task completion.
6.2 Human evaluation on RealEdit test set
Methods like VIEScore[29] align more closely with human judgment, but rely on vision-language models, which often miss subtle differences and produce inconsistent results.
Table 3: Elo rankings on RealEdit and GenAI Arena [23] test sets. On RealEdit test set, our model scores the highest. We perform competitively on GenAI Arena test set of synthetic data.
Model RealEdit GenAI Arena
Elo 95% CI Elo 95% CI
AURORA [28] 1019 +14/-11 - -
HIVE [66] 997 +16/-10 - -
InstructPix2Pix [6] 984 +13/-16 1011 -50/+47
MagicBrush [65] 982 +11/-13 1107 -39/+47
Null-text Inv. [36] 949 +10/-11 - -
SDEdit [35] 885 +13/-13 991 -48/+35
RealEdit (Ours) 1184 +17/-12 1043 -12/+17
To counteract this, we conducted a qualitative evaluation using Elo scores, following the methodology from GenAI Arena [23] and LMSYS [70]. This evaluation, conducted via Amazon Mechanical Turk, involved pairwise comparisons against the baselines on 200 diverse images from our test set. Results in Table 3 demonstrate that our model outperforms baselines on human judgement.
6.3 Deploying our model on Reddit
One limitation of standard Elo evaluations is that they are conducted by individuals with no personal connection to the image. To ecologically validate the utility of our model with photo owners, we deploy our model back on Reddit. We provide editing services for new user requests, posting edited images in the comments per subreddit guidelines.
On multiple occasions, we received positive feedback. For example, the model successfully removed red-eye from a photo. The original poster (OP) responded with: “Thank you so much! Solved.” and closed the request. On another occasion, we edited a picture of a car, and the OP remarked, “It looks pretty good, man.” On an edit of removing a person from the background, OP commented “Wow this looks great! I love the way you smoothed out the lighting on me as well” indicating that our model not only is successful semantically but produces aesthetic images.
Figure 5: Real requests completed on Reddit. We deployed our model on r/PhotoshopRequest to complete in-the-wild requests. We received positive feedback from users on the examples above.
6.4 Evaluations on existing test sets
We also conduct evaluations on external test sets including the test sets in GenAI Arena [23], Emu Edit [48], and MagicBrush [65]. On GenAI Arena, we report Elo ranking in Table 3 computed with real human preferences. Our model ranks second among the evaluated models. While these results were insightful, we found the examples in GenAI Arena to be less representative of real-world tasks. We include full automated evaluation results on Emu Edit and MagicBrush in Appendix, where our model performs competitively with the individual strongest models on respective test sets across varying metrics.
6.5 Improving edited image detection
Table 4: Model performance on detecting edited images in the RealEdit test set and in-the-wild images on <REDACTED>.
RealEdit dataset In-the-Wild dataset
[…]
Baseline 23.5 14 80 49 35 80
Ours 69 64 74 63 57 71
Change + 45.5 +50 -6 +14 +22 -9
Figure 6: The baseline misclassifies both images as real, whereas our model correctly spots the fake (right) that spawned the 2005 Paris Hilton “Stop Being Poor” meme.
We partnered with <REDACTED>, a platform where users can upload media to assess authenticity. Their primary fake image detection model is a fine-tuned version of Universal Fake Detect (UFD)  [38], which effectively detects model-generated deepfakes. We leverage the human-edited images in RealEdit to enhance the model’s ability to detect such edits, which has significant real-world impact.
UFD is trained on a recipe of 62K images from academic datasets [56, 25, 26, 24, 31, 51] and some proprietary data, none of which includes human edits. We trained a model from scratch using the UFD training pipeline with added RealEdit training data. We evaluated on the RealEdit test set and on a random subset of 100 reals plus 100 in-the-wild edited images from <REDACTED>. We show in Table 4 that fine-tuning on RealEdit improves F1 by 45.5 and 14 points on RealEdit and <REDACTED>’s test sets respectively. RealEdit also serves as a challenging human-edited image detection benchmark for models that are more specialized for this task compared to deepfake detection [52, 67, 12, 68].
Figure 7: Examples of the RealEdit model on RealEdit test set images compared to other editing models. Our edits are often more semantically correct as well as more visually appealing.
7 Discussion
Privacy and ethics.
To protect user privacy, individuals can opt out of having their images in the dataset by removing the photos from Reddit. Since our dataset contains image URLs rather than image files, images deleted from the web are automatically removed. Additionally, we provide a form where individuals can request their data to be removed from the dataset. Although this evolving dataset may introduce challenges for quantitative validation, ensuring user privacy remains our top priority.
Our work has positive social impacts, such as reducing the need for professional editing software and skills, and enabling higher-quality restorations of family photographs. However, we recognize the risks of malicious exploitation and strongly oppose any harmful, offensive, or derogatory use of our model or data. We plan to further pursue the development of fake image detection tools.
Conclusion.
We propose RealEdit, a dataset of 57K input - instruction - outputs data points where all instructions and edits are performed by humans. We analyze the distribution of real-world editing requests and fine-tune InstructPix2Pix [6] to create a SOTA image editing model on these tasks. Lastly we explore RealEdit ’s potential in facilitating deepfake detection.
Supplementary Material
Appendix A Data taxonomy
A.1 Full taxonomy
We include the taxonomies of RealEdit (Figure 8), Emu Edit (Figure 9), and MagicBrush (Figure 10) test sets, as well as the unabridged comparison between all three (Figure 11). The prompt used to taxonomize these requests is included in Figure 12. We notice RealEdit has a more diverse set of tasks as well as a more even distribution with greater focus in tasks like “remove” and “enhance”. Emu Edit [48] has a fairly even task distribution, though a smaller set of common tasks. MagicBrush [65] has a very skewed distribution, with a high focus on “add” tasks which are not likely to be requested by human users, as humans generally include all desired elements when taking a photograph.
Figure 8: Taxonomy of RealEdit image edit requests. There is a wide variety of task types and edit subjects, with subtle tasks like “remove” and “enhance” being the most requested.
Figure 9: Taxonomy of Emu Edit image edit requests. There is a smaller range of task types than RealEdit, but the distribution is fairly even.
Figure 10: Taxonomy of MagicBrush image edit requests. There is a limited selection and extremely uneven distribution of task types, with “add” accounting for almost half of all requests.
Figure 11: Differences in the distribution of our test set compared to MagicBrush and Emu Edit test sets. MagicBrush and Emu Edit tend to be similar in distribution to each other, but starkly different from RealEdit.
Figure 12: Prompt used for taxonomizing edit requests. We passed this along with input images to GPT-4o.
A.2 Performance across edit operations
We show the VIEScore comparisons of RealEdit, AURORA [28], InstructPix2Pix [6] and MagicBrush [65] in Table 5. We notice that in all of the editing tasks, the RealEdit model has the highest overall VIEScore. However, in “add” tasks, which comprise a much smaller percentage of our dataset compared to InstructPix2Pix and MagicBrush, we have a lower perceived quality, indicating that having more “add” data might improve the aesthetics. The task with the highest score for RealEdit is “remove”, with a VIE_O score of 4.35. The “remove” task comprises the largest portion of our dataset, which may explain this result. The hardest task is “formatting”, the only operation for which we do not have the highest semantic completion score. This is due to the fact that this task is impossible for current models to fulfill properly, as changing file formats, resizing, etc. are not supported by current model architecture.
Table 5: Breakdown of model performance by operation. We find that our model is consistently best across all operations in VIE_O, and our strongest operation is “remove”. We use a sample of 2000 data points and take arithmetic mean of all individual scores on each data point.
Operation AURORA InstructPix2Pix MagicBrush RealEdit
VIE_SC VIE_PQ VIE_O VIE_SC VIE_PQ VIE_O VIE_SC VIE_PQ VIE_O VIE_SC VIE_PQ VIE_O
Add 2.89 3.45 2.34 2.48 3.60 2.15 1.94 4.43 1.79 4.24 3.26 3.15
Change color 2.38 3.77 2.26 2.90 3.61 2.57 1.95 4.05 1.83 5.36 4.05 4.11
Enhance 1.86 3.00 1.88 1.80 2.91 1.79 2.41 3.44 2.33 4.73 4.03 3.95
Formatting 0.89 3.02 0.99 1.70 3.13 1.31 0.74 3.57 0.94 1.66 4.47 1.66
Open ended 2.51 2.70 1.98 2.49 3.57 2.05 2.36 3.49 1.99 4.67 2.93 3.15
Remove 2.94 4.25 2.76 1.01 3.03 1.06 2.30 4.71 2.30 5.29 5.01 4.35
Replace 2.18 3.32 1.87 2.25 3.16 1.74 1.57 3.81 1.45 3.50 3.50 2.53
Restore 1.52 2.23 1.57 1.66 2.49 1.74 1.59 2.60 1.74 4.01 2.98 3.21
Appendix B Data processing
Test set image captioning
We caption all input and ground truth images in the test set to enable evaluations with models that require captions. The process involves two main stages. First, for input image captioning, we pass the processed instruction along with the input image to LLaVA-Next[33]. This generates a caption for the input image that integrates the instruction, emphasizing key aspects of the image relevant to the editing task.
For output image captioning, we pass the input caption and edit instruction to GPT-4o, which combines these elements to generate a caption for the ground truth (edited) image, reflecting both the original content and the changes made according to the instruction. Refer to Figure 13 for examples of captions.
Figure 13: Examples of test set data with captions for input image and ground truth image.
Appendix C Discussion
C.1 Limitations and future work
RealEdit is collected from Reddit posts from 2012-2021. As such, we have less data and a danger of it getting outdated. We plan to regularly update our dataset to ensure that the edits reflect as current culture as much as possible. This will also help in edited image detection, by facilitating the detection of edits where newer AI tools were used, as the line between human editing and model editing is increasingly blurred.
We also filter our dataset in order to more closely match the training distribution, removing some natural diversity of human edit requests. In future work, we hope to explore different architectures capable of handling real world edit requests and editing styles.
The pretraining of the RealEdit model uses CLIP embeddings, which while very useful for semantic changes to an image, a large portion of the RealEdit dataset involves edits that do not involve semantic changes. Additionally, in edited image detection, some of the edits may not change the embeddings much. We urge future work to explore alternatives to such embeddings that may capture purely aesthetic changes.
C.2 Social impacts
The social impact of our dataset stems from both the effect on model training as well as the ability of our test set to be used to accurately and justly benchmark other models. The training data will inform how well the RealEdit model performs certain types of edits. The test set on the other hand determines the factors we incentivize in other models.
Accessible image editing models that are capable of handling real world tasks are extremely useful in democratizing the documentation of people’s lives. For example, some requests in RealEdit involve restoring old photographs, many of which were paid. The RealEdit model can help more users to document meaningful family histories, even if they cannot afford to pay for edits. We have demonstrated the efficacy of our model on making real world edits by uploading our model’s generations to Reddit. Additionally, our exploration of the contribution of RealEdit in deepfake image detection has shown that RealEdit increases the ability of <REDACTED>’s ability to detect fake images, which is extremely useful in a world where images are routinely edited to cause scandals or spread misinformation.
There is a known issue in image generation models of generating images or making edits based on demographic biases such as smoothing wrinkles, lightening skin, and male bias in certain professions, which may offend users. Additionally, our dataset mirrors the demographic profile of Reddit users, who are predominantly Western, younger, male, and left-leaning, potentially influencing the types of images and editing requests included. We hope to study the effect of this extensively in RealEdit in future work.
There is also an issue of inappropriate edits, which we have mitigated to our knowledge in RealEdit through filtering of NSFW content using opennsfw [4], along with manual filtering in our test set.
C.3 Ethics
Some other editing datasets [65] do not use human faces in order to evade biases as well as privacy concerns. However, in RealEdit, we determine that since over half of edit requests contain images focused on people, we must train on human data in order to be successful in completing real world editing tasks. To mitigate privacy concerns, we use the URL in place of the actual input image so that if the original poster (OP) deletes their post, it will be removed from our dataset. We also include a form for users to request their data to be removed. In the case of mitigating biases, we hope in future work to study the effects of using Reddit data on task completion for a wide array of demographic groups, as well as techniques or supplementary data sources to boost performance on underrepresented groups. This is a known problem in the field, and we are compelled by user preferences to include human data. Given this, although we appreciate the importance of mitigating demographic biases, this is outside the scope of a single paper.
Appendix D Modeling ablations
D.1 Implementation details
We fine-tune the checkpoint of InstructPix2Pix [6] using the RealEdit training set for 51 epochs on a single 80GB NVIDIA A100 GPU. The total batch size is 128, and the learning rate starts at $2\times 10^{-4}$ We resize images to 256 × 256, disable symmetrical flipping to maintain structural integrity, and apply a cosine learning rate decay to $10^{-6}$ over 15,000 steps with 400 warmup steps. The training process takes 24 hours.
D.2 Consistency decoder
Table 6: The decoder has minor effects on quantitative metrics but sometimes improves qualitative results.
Model VIE_O VIE_PQ VIE_SC L1 L2 CLIP-I DINO-I CLIP-T
RealEdit w/ original decoder 3.54 3.91 4.37 0.154 0.069 0.830 0.782 0.258
RealEdit w/ consistency decoder 3.48 3.78 4.34 0.156 0.069 0.830 0.779 0.258
Change -0.06 -0.13 -0.03 0.002 0 0 -0.003 0
MagicBrush w/ original decoder 1.92 3.98 1.89 0.139 0.066 0.830 0.782 0.251
MagicBrush w/ consistency decoder 1.84 3.93 1.83 0.135 0.066 0.831 0.784 0.251
Change -0.08 -0.05 -0.06 -0.004 0 0.001 0.002 0
InstructPix2Pix w/ original decoder 1.73 3.37 1.85 0.183 0.075 0.754 0.651 0.243
InstructPix2Pix w/ consistency decoder 1.89 3.40 1.95 0.180 0.073 0.758 0.648 0.244
Change 0.16 0.03 0.10 -0.003 -0.002 0.004 -0.003 0.001
We integrate OpenAI’s Consistency Decoder [40], which is designed to enhance the quality of specific features during inference. This has a minimal impact on overall model performance metrics but proves highly effective for improving the handling of faces, textures, and intricate patterns.
As the decoder operates independently of the underlying model, we evaluate its effectiveness with InstructPix2Pix[6] and MagicBrush[65] on a sample of 500 tasks. The results indicate that while the decoder minimally affects standard metrics, such as VIEScore[29] and CLIP-T (Table 6), it often enhances the aesthetic quality in areas requiring fine detail, such as facial reconstruction and complex textures (Figure 14).
These findings demonstrate the decoder’s potential as a lightweight, inference-only addition to improve the output quality of existing image-editing models without altering their core architectures or diffusion processes.
Figure 14: Consistency decoder allows for more aesthetic generation of faces.
D.3 Data filtering
We observe that human-generated edits often introduce substantial diversity, such as rearranging objects or people, which significantly impacts Structural Similarity Index Measure (SSIM) scores. These variations create a distributional mismatch with InstructPix2Pix’s pretraining data (Figures 15, 16, 17), where edits are generally more constrained. To better understand this difference, we analyze SSIM distributions, highlighting the gap between human edits and the structured outputs of synthetic datasets.
To make our dataset more compatible with InstructPix2Pix, we currently apply SSIM-based filtering to exclude edits that deviate too far from the pretraining distribution. Following this, we use the same CLIP-based filtering methodology employed by InstructPix2Pix to further refine the data. We verify that this filtering leads to a more capable model using the VIE-scores (Table 7) and CLIP-based metrics (Figure 18). Our approach relies on thresholding to identify and remove outliers, but we recognize that soft sampling techniques could offer a more flexible and nuanced alternative. Exploring such methods remains a promising direction for future work.
Table 7: Aligning RealEdit data to the pretraining distribution yields better results.
Model VIE_O VIE_PQ VIE_SC
Filtered data 3.48 3.78 4.34
Original data 2.35 2.99 2.91
D.4 Processing instructions
Reddit users often provide vague, unclear instructions with unnecessary details, hindering the editing process. To address this, we refined these instructions for greater clarity and relevance. To evaluate the impact of this preprocessing, we trained two models under the same conditions: one with the original instructions and the other with the processed versions. Results in Table 8 and Figure 20 show that these have a significant effect on model performance.
We ran this experiment early in the development processes with a suboptimal training strategy and a smaller subset of the data, leading to much lower scores compared to our final model.
Table 8: Processing instructions improves model performance.
Model VIE_O VIE_PQ VIE_SC
Processed instructions 2.42 3.72 2.84
Original instructons 2.06 3.10 2.45
Figure 15: SSIM distribution of InstructPix2Pix training data.
Figure 16: SSIM distribution of MagicBrush training data.
Figure 17: SSIM distribution of RealEdit training data.
Figure 18: Filtering the data massively improved CLIP-based metrics.
Figure 19: GPT-4o prompt for instruction rewriting.
Figure 20: Processing instructions consistently yields better results on CLIP-based results.
Appendix E Inference time results
E.1 Hyperparameters
We conducted several inference-time experiments: varying the number of diffusion steps, the image and text guidance scales, and further rewriting instructions with GPT-4o to add more details.
See equation (6) in  [21] for the definition of classifer-free guidance scale. The conventional wisdom is that higher image guidance scale make the generated image look more similar to the original image, while higher text guidance scale improve instruction adherence. Additionally, higher number of inference steps are believed to improve the quality of the generated image at the expense of computational time. Our statistical experiments do not capture these relationships, and even demonstrate the opposite relationship in case of image guidance scale.
Number of inference steps
We observe that 20 inference steps strike a good balance between the computational time and the image quality. Specifically, we find that the average CLIP similarity between the generated image and the most upvoted Reddit edit is approximately the same for any setting of inference steps above 20. See Figure 22 for the statistical plot and figure 21 for an example.
Figure 21: Increasing the number of diffusion steps above 20 usually does not improve the quality.
Figure 22: The number of inference steps does not improve the generated image quality, as measured by the CLIP similarity between the generated image and the most upvoted Reddit edit.
Text guidance scale
[…]
Figure 23: An example where guidance scales behave as expected.
Figure 24: Text guidance scale has no effect on instruction adherence, as measured by the CLIP similarity between the generated image and the caption of the expected output, as in figure 13.
Image guidance scale
[…]
Figure 25: Increased image guidance scale results in less similar images, as measured by CLIP similarity between the input and generated images.
E.2 Instruction rewriting
As the diffusion model lacks reasoning capabilities, it often fails when asked to interpret abstract or creative instructions. To improve outcomes on these examples, we employ a large language model (LLM) to rewrite instructions in a more specific manner, similar to Dalle-3  [3]. Since only creative edit tasks benefit from this technique, we do not make this part of our main pipeline. We gave the input image and the original instruction to GPT-4o with the prompt “You are given an image editing instruction. If the instruction is already concrete and specific, do not rewrite it at all. If the instruction is vague or does not make sense for the image, then rewrite it. Make the new instruction specific and detailed, e.g. do not use words ’enhance’, ’adjust’, ’any’.
Figure 26: Detailed instructions can improve edit quality on certain classes of tasks.
E.3 Quantitative evaluation on external test sets
Table 9: Evaluation on MagicBrush and Emu Edit test sets. All scores within 1 standard deviation of the highest score are underlined. The RealEdit model is still able to perform competitively on some metrics despite these tasks being out of distribution.
MagicBrush Test Set Emu Edit Test Set
Model VIES_SC $\uparrow$ VIEPQ $\uparrow$ VIEO $\uparrow$ VQA_llava $\uparrow$ VQA_CLIP $\uparrow$ TIFA $\uparrow$ VIES_SC $\uparrow$ VIE_PQ $\uparrow$ VIE_O $\uparrow$ VQA_llava $\uparrow$ VQA_Flan-t5 $\uparrow$ TIFA $\uparrow$
AURORA [28] 4.11 3.86 5.52 0.5179 0.6517 0.6968 3.40 4.86 3.81 0.4923 0.6178 0.6705
Emu Edit [48] N/A N/A N/A N/A N/A N/A 4.66 5.11 5.72 0.5130 0.6489 0.6692
HIVE [66] 2.86 5.02 3.43 0.5200 0.6547 0.6918 1.89 5.50 2.06 0.4372 0.5258 0.6447
InstructPix2Pix [6] 2.63 4.70 3.06 0.4490 0.5518 0.6615 2.15 5.00 2.36 0.4261 0.5061 0.6343
MagicBrush [65] 3.43 4.89 4.11 0.5554 0.7138 0.7103 2.91 5.47 3.13 0.4680 0.5808 0.6628
Null-text Inv. [36] 2.77 4.74 3.29 0.5246 0.6429 0.6899 3.43 5.10 3.93 0.4823 0.5931 0.6578
SDEdit [35] 0.90 2.26 1.02 0.4185 0.4191 0.6167 0.95 3.23 1.06 0.4406 0.5145 0.6417
RealEdit 3.12 3.60 4.09 0.5088 0.6299 0.6865 3.27 4.86 3.84 0.4938 0.6158 0.6650
Despite being out of distribution, the RealEdit model performs comparably to other models on the synthetic datasets Emu Edit [48] and MagicBrush [65]. On several metrics (VQA_CLIP and TIFA on MagicBrush and VQA_llava, VQA_Flan-t5 and TIFA on Emu Edit), the RealEdit model is within 1 standard deviation of the highest scoring model, indicating that it is fairly generalizable to new tasks.
E.4 Elo scores
To evaluate Elo scores, we leverage Amazon Mechanical Turk (MTurk) for conducting pairwise comparisons. We selected 200 diverse examples from our dataset to ensure coverage of various editing tasks and performed comparisons across all seven models in our benchmark. This process resulted in a total of 4,200 pairwise evaluations, providing a robust dataset for assessing human preferences. We present a table of pairwise winrates (Figure 27)
In addition to evaluating our dataset, we extended our analysis to the Imagen Hub Museum[30] tasks, building on the results from the GenAI Arena[23]. Using their generations, available on HuggingFace, we incorporated results from our model to facilitate direct comparisons. For these evaluations, we conducted a new round of pairwise comparisons where we matched one model from their benchmark against our model for the same tasks. This allowed us to directly assess how our model performs relative to state-of-the-art models on external datasets.
The evaluations on MTurk followed a structured protocol to ensure reliability and consistency. Workers were asked to compare image outputs based on task completion, realism, and alignment with instructions. The use of MTurk enabled us to gather diverse human feedback efficiently and at scale. The full results are presented in Table 10, highlighting the comparative performance across different models.
Table 10: Elo scores of models based on the GenAI[23] test set.
Model Elo Rating 95% C.I. Sample Size
MagicBrush [65] 1107 -39/+47 132
CosXLEdit [1] 1064 -49/+42 133
RealEdit 1043 -12/+17 1117
InfEdit [59] 1023 -44/+39 122
InstructPix2Pix [6] 1011 -50/+47 117
Prompt2prompt [20] 1011 -46/+46 119
PNP [53] 992 -43/+62 122
SDEdit [35] 991 -48/+35 126
CycleDiffusion [57] 933 -41/+49 120
Pix2PixZero [42] 834 -46/+41 126
Figure 27: Heatmap of pairwise winrates on our test set. We excluded draws for this heatmap.
Figure 28: Interface for Elo evaluation on MTurk. To complete Elo evaluations, we hired workers on Amazon Mechanical Turk to compare the quality of different editing models.
Appendix F Reddit experiment
To evaluate the generalization capability of our model, we deployed it on Reddit. Specifically, we targeted two subreddits: r/PhotoshopRequest and r/estoration, which focus on image editing and restoration tasks. Adhering to the community guidelines of these subreddits, we collected posts requesting image edits and processed them using our model.
For each processed request, we submitted a comment containing the generated output image along with a brief message asking for feedback from the original poster (OP). With this experiment, we gathered qualitative evaluations from humans, and provide insight into the model’s performance in real world scenarios. See Figures 29 and 30.
Figure 29: Our model successfully completes new requests on Reddit. Deployed on the original subreddits, it handled in-the-wild requests effectively as seen by OP’s response.
Figure 30: Our model successfully completes new requests on Reddit. Deployed on the original subreddits, it handled in-the-wild requests effectively as seen by OP’s response.
Appendix G Edited image detection
Data processing and training
The baseline classifier undergoes a multi-stage training process: initially on academic datasets and subsequently fine-tuned on <REDACTED>’s proprietary data. In total, the baseline model is trained on 65K images with a near equal 50/50 split between real and generated images. To assess the value of RealEdit data for fake image detection, we train a second version of UFD by combining the original data with RealEdit data. Specifically, we include only photographs, excluding non-photographic images such as digital artworks, screenshots, cartoons, and infographics, filtered using GPT-4o. This single-stage training incorporates an additional 37K original and 37K edited images, resulting in a total of 139K images.
In the first stage of training, the <REDACTED> model took over 24 hours to train on an A10G GPU with 20GB of RAM and the remaining three stages took 4 hours. Our optimized model took 1.5 hours to train on a L40S GPU with 40GB of RAM.
Table 11: Breakdown of fake image sources in the training recipe of the <REDACTED> model used as our baseline.
Source Count
DiffusionDB  [56] 16K
StyleGAN2-FFHQ  [26] 8K
Stable-Diffusion-Face  [51] (512 resolution) 2.4K
Stable-Diffusion-Face (768 resolution) 2.4K
Stable-Diffusion-Face (1024 resolution) 2.4K
Fakes uploaded to <REDACTED> 2K
Table 12: Breakdown of real image sources in the training recipe of the baseline model.
Source Count
CelebA-HQ (Reals)  [24] 23K
Random sample of COCO-Train-2017  [31] 5K
Flickr-Faces-HQ Dataset (FFHQ)  [25] 3K
Reals uploaded to <REDACTED> 0.7K
<REDACTED>’s in-the-wild test set
<REDACTED>’s in-the-wild test set includes images uploaded between 8/16/2024 and 11/10/2024. We randomly sample 100 real images and then sample 100 fake images selected from those tagged as ”likely photoshopped” by professional sleuths in <REDACTED>’s media database, ensuring the evaluation focuses on human-edited images rather than exclusively AI-generated edits. Tool usage data was available for some images, revealing that approximately 80% of the fake images were human edits created with Photoshop, while the remaining 20% involved human edits combined with AI tools such as Dream Studio AI, Insightface AI, and Remaker AI.
Qualitative example
To understand how the classifier operates, we use Gradient-weighted Class Activation Mapping (Grad-CAM) [47] to analyze an example. In Figure 31, we show an edited image where a bear was added to the background using Photoshop. The original image did not include the bear. The baseline model incorrectly classified this photo as unedited, whereas the classifier trained with RealEdit data correctly identified it as edited. Grad-CAM highlights the areas of the image most influential to the classifier’s decision, as seen in the figure, where the focus is on the region around the bear. The specific implementation we adapted is from Gildenblat and contributors [17].
Figure 31: Top: An edited image that inserted a bear to make it seem the camera crew was being chased. Bottom: Grad-CAM heat-map visualization highlighting the regions of attention.
Appendix H Additional results
Figure 32: Additional examples of RealEdit generations on RealEdit test set compared to all other baseline models. We notice that the RealEdit model consistently outperforms other models in task completion as well as aesthetic quality.
Figure 33: Additional examples of RealEdit generations on RealEdit test set compared to select high performing baseline models. We notice that the RealEdit model consistently outperforms other models in task completion as well as aesthetic quality.